AddyM:

hey can someone help me I don't need the answer I just need an example on how to do it. will post screen shot in a sec

AddyM:

1 attachment
AddyM:

i forgot how to do it :p

snowflake0531:

do you still need help?

AddyM:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 do you still need help? \(\color{#0cbb34}{\text{End of Quote}}\) yes. sorry I was @ gym

AddyM:

help also @RomMosher go away

snowflake0531:

do you know how to use the substitution method?

snowflake0531:

or not at all

snowflake0531:

the elimination method is kind of like subtracting or adding whole equations for example in this one -8x+5y = 2 - -12x + 5x = -2 4x = 4 x=1

snowflake0531:

now that we have x, we can substitute x into the equation, both works i'll just do both to show you

snowflake0531:

-8x+5y = 2 -8(1) + 5y =2 5x =10 x=2

snowflake0531:

-12x + 5y = -2 -12(1) + 5y -2 5y =10 y=2

snowflake0531:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 the elimination method is kind of like subtracting or adding whole equations for example in this one -8x+5y = 2 - -12x + 5x = -2 4x = 4 x=1 \(\color{#0cbb34}{\text{End of Quote}}\) i meant -12x+5y

snowflake0531:

so, the point of intersection is (1,2)

AddyM:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 so, the point of intersection is (1,2) \(\color{#0cbb34}{\text{End of Quote}}\) jesus just saw this sorry thank you! :p

snowflake0531:

yw~